9. (a) Show that the set {( , 3|) 5 15}

2 2

S = x y x + y ≤ is convex.


1
Expert's answer
2021-05-03T05:57:17-0400

Let C be the convex set,


"C = { x\in R^2 | u.v\le9}"


The set C is convex if for any "u,w \in C" the point


"tu+(1-t)w \in C" for all "t\in[0,1]."


Let "u,w \in C" and "t \in [0,1]."


"(tu+(1-t)w).v"


"= (tu).v+[(1-t).w].v"


"= tu.v+(1-t)w.v"


"\le t\times 9 + (1-t)\times 9"


"= 9"


Therefore, "(tu+(1-t)w).v \le 9"


Hence, C is convex.